OpenStudy (anonymous):

Which shows 51^2 - 49^2 being evaluated using the difference of squares method?

OpenStudy (anonymous):

512 - 492 = 2601 - 2401 = 200 512 - 492 = (51 + 49)(51 - 49) = (100)(2) = 200 512 - 492 = (51 - 49)2 = 22 = 4 512 - 492 = (2601 + 2401)(2601 - 2401) = 1,000,400

OpenStudy (anonymous):

C. because difference of squares is like factoring. Taking two numbers and then squaring or factoring them (depending if it's a quadratic equation or not).

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

Ya sure!

OpenStudy (anonymous):

Which of the following best describes the graph of f(x) = x2 - 3x - 10?

OpenStudy (anonymous):

Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (jdoe0001):

easier if you post anew, more eyes

OpenStudy (anonymous):

That would be A. Look at www.desmos.com the calculator will be useful. Just type the equation and then see what points land on that equation.

OpenStudy (anonymous):

Thanks :D I wish I could give you two medals Dx

OpenStudy (anonymous):

Lol you can give me a medal for this question I answered for someone else: http://openstudy.com/study#/updates/55c29d5ae4b08de2ded201be

OpenStudy (anonymous):

And you're welcome! :D